### v3

# DeCoF: Generated Video Detection via Frame Consistency: The First Benchmark Dataset

Abstract
The escalating quality of video generated by advanced video generation methods results in new security challenges, while there have been few relevant research efforts: 1) There is no open-source dataset for generated video detection, 2) No generated video detection method has been proposed so far. To this end, we propose an open-source dataset and a detection method for generated video for the first time. First, we propose a scalable dataset consisting of 964 prompts, covering various forgery targets, scenes, behaviors, and actions, as well as various generation models with different architectures and generation methods, including the most popular commercial models like OpenAI’s Sora and Google’s Veo. Second, we found via probing experiments that spatial artifact-based detectors lack generalizability. Hence, we propose a simple yet effective detection model based on frame consistency (DeCoF), which focuses on temporal artifacts by eliminating the impact of spatial artifacts during feature learning. Extensive experiments demonstrate the efficacy of DeCoF in detecting videos generated by unseen video generation models and confirm its powerful generalizability across several commercially proprietary models. Our code and dataset will be released at https://anonymous.4open.science/r/DeCoF-8394.
1 Introduction
Following the success of diffusion models in generating highly realistic images [28, 34, 30], increasing attention has been paid to video generation [21, 4, 51], enabling the creation of high-quality videos from general conditional inputs such as text and images. However, this powerful generation capacity also introduces a new crisis of trust [36, 3]. Therefore, there is an urgent need for detectors for generated videos.
Figure 1: Can the detector trained on a specific video generation model detect videos generated by unseen generation models?
In this paper, we tackle an overlooked challenge, i.e., generated video detection, by proposing a universal detector to deal with the trust crisis caused by various emerging text-to-video (T2V) models, especially unseen models. Moreover, we propose the first comprehensive generated video detection dataset, the $GeneratedVideoForensics$ (GVF) dataset, for evaluating the performances of generated video detectors. GVF encompasses a wide range of scene contents and motion variations, approximating the authenticity verification problem caused by the T2V generation model in diverse actual environments. Pairs of real and generated videos in the dataset are constructed using the same prompt to ensure that the generated video detectors learn subtle distinctions to identify the authenticity of videos rather than learning differences between the content of the real and fake videos. Furthermore, we construct generated samples corresponding to real videos based on various T2V models with different generation methods and model architectures, such as pixel-space or latent-space models [22, 51, 46, 42], to measure the generalizability of the generated video detectors.
(a) Geometry
(b) Layout
(c) Frequency
(d) Color
(e) Temporal
Figure 2: Illustration of spatial and temporal artifacts on generated video. Spatial artifacts: (a) errors in geometric appearance, (b) errors in image layout, (c) frequency inconsistency, (d) color mismatch; Temporal artifacts: (e) mismatch between frames.
Considering that authentic videos should present smooth inter-frame consistency, a naive approach is to utilize end-to-end spatiotemporal convolutional neural networks (STCNN) for generated video detection. However, we found via exploratory experiments that these models lack generalizability to detect videos generated by unseen T2V models.
Further probe experiments revealed that spatiotemporal convolutional models primarily recognize video authenticity through spatial artifacts (Figure 2 illustrates two types of artifacts in generated videos), degrading video detection into a 2D problem. By analyzing the performance of existing text-to-image (T2I) detection works [48, 43, 47], we found it difficult to capture general spatial artifacts, as the contrast and style of video frames generated by different generation models vary.
Therefore, to develop a universal generated video detector, we propose the first temporal-artifact-based detection method. More specifically, we decouple temporal and spatial artifacts by mapping video frames to feature space where the inter-feature distance is negatively correlated with image similarity, to detect the anomalies caused by inter-frame inconsistency. Our method greatly reduces computational complexity and memory requirements since only anomalies between features need to be learned. Extensive experiments demonstrate that our method exhibits extraordinary capabilities in generated video detection and excellent generalizability for unseen generation models, even on the latest generation models Sora [41] and Veo [45].
Our contributions can be summarized as follows:
  • We present the first comprehensive and scalable dataset for benchmarking generated video detectors.
  • Through a series of probing experiments, we have demonstrated the flaws in detecting generated videos from spatial artifacts.
  • We develop the first temporal-artifact-based universal generated video detector named DeCof and validate its performance through extensive experiments, including tests on the latest commercial proprietary models.
2 Related work
2.1 Generation Models
Recently, with the success of diffusion models in generating images [20, 40, 28, 9, 34, 31], a number of research explores text-to-video Diffusion Models (VDMs). Some works [33, 4, 22, 46] ground their models on latent-based VDMs, e.g., Text2Video-zero [22] generates videos by using motion dynamics to enhance the latent code of text-to-image(T2I) model [34] without the need for training on any video data, and ModelScopeT2V [46] extends T2I model with temporal convolution and attention blocks trained with image-text and video-text datasets. ZeroScope [42] is a watermark free version of ModelScopeT2V, and has been optimized for 16:9 composition. On the other hand, PYoCo [14] and Make-A-Video [38] anchor their models on pixel-based VDMs. Show1 [51] integrates Pixel-based VDMs to produce videos of lower resolution with better text-video alignment, then upscale them with latent-based VDMs to create high-resolution videos with low computation cost.
T2V-zero
Show1
MScope
ZScope
Real
Time 0
Time 1
Time 2
Time 3
Time 4
Figure 3: We generated four sub-datasets based on the real videos using the prompts and selected generation models employing four different architectures and generation methods: Text2Video-zero, ModelScopeT2V, ZeroScope, and Show-1. Each pair of positive and negative samples use the same prompt. Here shows the first five frames of four sample pairs using the prompt: A person is cutting a carrot.
2.2 Generated Content Detection
Many studies have focused on generated image detection [27, 52, 2, 7, 35], especially for unseen generation models. A simple classifier trained by Wang et al. [47] on ProGAN-generated images can be generalized to 20 unseen other GAN-generated images. Frank et al. [13] conclude that the artifacts of the generated images are caused by the upsampling method. Wang et al. [48] identify generated images through the difference of images reconstructed by a pre-trained diffusion model. Ojha et al. [29] propose to develop a universal generated image detector based on pre-trained vision-to-language models. Meanwhile, a large amount of work focuses on Deepfake image detection [16, 23, 53, 37], using methods such as artificially creating general forged boundaries [37], detecting shifts in the frequency domain [18], and detecting inconsistencies in face image regions [54]. So far, there are some studies on detecting Deepfake videos [17, 26, 1, 19, 49], but research on detecting generated videos remain unexplored. This paper targets at bringing exploratory and leverageable contributions to this new direction.
3 Problem Formulation
Generated Video Detection. In addition to the generalizability to unseen generation models, we also take the generalizability to cross-category videos into account. The generated video detection task can be defined as follows:
Given a collection of real videos denoted as $R=\{r_{1},r_{2},\cdots,r_{n}\}$, a set of content prompts $P=\{p_{1},p_{2},\cdots,p_{n}\}$, where $r_{i}$ and $p_{i}$, $i\in[1,n]$ refer to each real video and its corresponding content prompt. Let $G=\{g_{1},g_{2},\cdots,g_{m}\}$ denote a set of video generation models. Using the prompt $p_{i}$ of each real video $r_{i}$, $G$ generates $m$ corresponding videos, resulting in a set of fake video collections $F=\{F_{1},F_{2},\cdots,F_{m}\}$, where $F_{j}=\{g_{j}(p_{1}),g_{j}(p_{2}),\cdots,g_{j}(p_{n})\}$, $j\in[1,m]$. As such, we form the generated video detection dataset $\mathcal{D}=\{D_{1},D_{2},\cdots,D_{m}\}$, where $D_{j}=\{(x,y)^{(i)}\}_{i=1}^{n}$, with $\mathcal{X}=\{R,F_{j}\}\subset\mathbb{R}^{L*H*W*C}$, $\mathcal{Y}=\{0,1\}$, $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ denoting a video and its label, $L,H,W$ and $C$ correspond to the video length, video frame height, width and number of channels, respectively.
The goal is to train a detector $\mathcal{F}$ on a subdataset $D_{j}$ built based on a known generation model $g_{j}$, map the video space to the category space $\mathcal{F}$ :::: $\mathcal{X}\rightarrow\mathcal{Y}$, and generalizes to unseen generation models $G_{k}$, where $k\in[1,m],k\neq j$. This can be achieved by minimizing classification error of $\mathcal{F}$ on training data $D^{*}_{i}$:
$\mathop{\arg\min}\limits_{\theta}\mathbb{E}_{(x,y)\in D^{*}_{i}}\mathcal{L}(f(x),y)$ (1)
where $\mathcal{L}$ is a loss function, and $\theta$ are the trainable parameters of detector $f$, $D^{*}_{i}$ refers to the training set of $D_{i}$. We evaluate the detection of cross-category generated videos by ensuring that the training set and the test set do not contain videos using the same prompt.
4 Dataset Construction
To benchmark generated video detectors, we have created the first dataset for the generated video detection task: the $GeneratedVideoForensics$ (GVF) dataset.
4.1 Dataset Composition
The GVF dataset consists of 964 triples and can be represented as $\mathcal{D}=\{(r,p,F)^{(i)}\}_{i=1}^{964}$, where $r$ and $p$ represent the real video and the corresponding text prompt, and $F$ represents the videos generated by the generation models using the text prompt $p$. In order to examine the generalizability of the detectors, we have selected four open source generation models employing different generation methods and model frameworks, namely Text2Video-zero [22], ModelScopeT2V [46], ZeroScope [42] and Show-1 [51] as generation models for negative samples. Accordingly, the GVF dataset could be divided into four sub-datasets. Figure 3 shows an example triple.
Figure 4: Data distribution across categories under the “major content (upper) and “attribute control" (lower) aspects.
Figure 5: Performance comparison of the 4 selected models across spatial major contents.
Prompt. To approximate the distribution of videos in reality,we have collected 964 prompts and corresponding real videos from open-domain text-to-video datasets: MSVD [6] and MSR-VTT [50]. Following FTEV [25], we consider the collection of prompts from the following four aspects to simulate the sample distribution in reality: the spatial and temporal content it mainly describes and the spatial and temporal attributes it controls during video generation. Spatial content is classified as people, animals, plants, food, vehicles, buildings, and artifacts, with corresponding spatial attributes of quantity, color, and camera view. Temporal content encompasses actions, kinetic motions, fluid motions, and light change, with attributes including motion direction, speed, and event order. This ensures that our GVF dataset is comprehensive and balanced. Figure 5 illustrates the data distribution across categories under different aspects, which indicates that the GVF dataset contains enough data to cover different category ranges.
4.2 Dataset Evaluation
Among many excellent T2V generation models, we have selected four of the most representative open-source models, including different generation methods and model architectures, to simulate the challenges posed by unseen generation models. Figure 5 shows the generation ability of the selected generation models on different attributes under the FETV [25] standard. Clearly, each model has its advantages and disadvantages. Text2Video-zero, for example, has the best performance on the spatial quality of the video, but the worst performance on the temporal continuity. On the contrary, ZeroScope has a better performance on temporal continuity but the worst performance on spatial quality. The choice of these video generation models ensures that our dataset can effectively measure the performance of detectors when facing unseen T2V generation models.
Table 1: Spatiotemporal Convolutional Neural Networks (STCNN) rely on spatial artifacts.
Method Artifacts Difference
Temporal Spatial
ACC AP ACC AP ACC AP
I3D 51.17 51.61 86.67 98.61 35.50 47.00
Slow 50.04 50.37 93.81 99.51 43.77 49.14
X3D 50.04 54.31 78.39 92.84 28.35 38.53
Mvit 50.48 55.29 72.16 92.59 21.68 37.30
5 Methodology
5.1 Probe experiments of STCNN models
Initially, we considered using spatiotemporal convolutional model(STCNN) [5, 12, 11, 10], which are widely used in video classification tasks, to train the generalizable generated video detector. However, we discovered that these models are only effective on the specific generation models they were trained on, failing to discern unseen T2V generation models.
To analyze what information the STCNN network was learning during the training process, or rather, which features within the video it utilized to distinguish between authentic and fabricated content, we designed a probing experiment with two novel datasets under the same model framework. The first dataset involves scrambling the frames of all real videos in the current test set, destroying the temporal continuity of the videos, to observe if the model can differentiate based solely on the temporal discontinuities of the videos. The second dataset is created by randomly selecting a frame of fake videos from the current test set and replicating it multiple times to observe if the model could differentiate based on the spatial artifacts of the video frame alone.
The experiment results in Table 1 show that STCNN overly relies on spatial artifacts, degrading the generated video detection into a 2D problem. However, the fact STCNN relies on spatial artifacts still does not explain its lack of generalizability to unseen generation models.
To further explore the possibility of detecting generated video based on spatial artifacts, we turned to the SoTA generated image detection works.
5.2 Capturing Universal Spatial Artifacts is Difficult
Figure 6: t-SNE visualization of real and fake video frames associated with four video generation models.
We transform the current detection of generated images into image level detectors for detecting generated videos and test them on the GVF dataset. As shown in Table 2, these detectors demonstrate superior capabilities in capturing the spatial artifacts of specific generation models. However, they also lack generalizability to unseen generation models.
To find out why these detectors lack generalizability to unseen generate models, we visualized the feature space of the detector [47] and examined the distribution of five sub-datasets: (i) Real, comprising real videos from the test set, (ii) T2V, comprising fake videos from the sub-dataset Text2Video-zero, (iii) Show1, comprising fake videos from the sub-dataset Show-1, and (iv) MP, comprising fake videos from the sub-dataset ModelScope, (v) ZP, comprising fake videos from the sub-dataset ZeroScope.
For each video, we utilized the detector trained on (i) and (ii) to extract their feature representations corresponding to each frame and subsequently visualized these through t-SNE [44].
Real
T2V
Show1
MP
ZP
Figure 7: Average spectrum of video frames for real videos and fake videos generated by different video generation models.
As shown in Figure 6, on the one hand, the detector has successfully classified the real video and the fake video generated by the seen generate model (Text2Video-zero), indicating that the detector has fully captured the spatial artifacts of the videos generated by the seen generate model. On the other hand, the detector cannot distinguish fake videos generated by unseen models from the real videos. Why are the spatial artifacts captured by the detector in (ii) not helpful for detecting videos generated by other video generation models? We illustrate this by visualizing the frequency spectrum of video frames generated by the models. As shown in Figure 7, although the generated videos all have spatial artifacts, such as parallel slashes appearing in T2V and bright spots appearing in Show-1, these spatial artifacts are not consistent. Therefore, capturing universal spatial artifacts appears to be extremely difficult.
Figure 8: Overview of the DeCoF framework. We first get real video and generated video features using the pre-trained CLIP:VIT, to eliminate the impact of spatial artifacts on capturing temporal artifacts. Then a verification module consisting of two transformer layers and one MLP head is used to learn the differences between frame consistency of the real and fake videos.
5.3 Detection based on Temporal Artifacts
The recent video generation models convert the T2I model into the T2V model through additional constraints, which have common drawbacks in controlling video continuity in two aspects: 1) the normativity of motion, not only do visible movements violate to physical laws, but more importantly, there are irregular movements of invisible detail pixels, 2) there is a slight pixel shift in the relatively static local background. Therefore, we propose to develop a detection method based on temporal artifacts.
Given a set of video frames $F=\{F_{1},F_{2},\cdots,F_{L}\}$, we define the method to determine the authenticity of a video based on its temporal artifacts as $f(F_{1},F_{2},\cdots,F_{L})=y$, where $f$ is a specific function that measures the consistency between video frames, $y\in Y=\{0,1\}$ is the authenticity label of the video. It appears that the detection model tends to learn spatial artifacts (Section 5.1). Therefore, we propose to eliminate the impact of spatial artifacts in order to help the model focus on capturing temporal artifacts.
Intuitively, we want to find a particular mapping function $M$ with the following properties: the distance between mapping frames is negatively related to the image similarity, and the mapping function is insensitive to spatial artifacts. Hence, we choose ViT-L/14 [8] trained for the task of image-language alignment as mapping function $M$. Then by mapping the video frames $F_{i}$ to $M(F_{i})$, We have implemented decoupling the temporal and spatial artifacts because the mapping function is not trained to capture spatial artifacts, but only encodes the image into a feature space with excellent properties, where the distance between features is inversely proportional to the image similarity. As a result, the original problem can be expressed as $f(M(F_{1}),M(F_{2}),\cdots,M(F_{L}))=y$.
Model framework. In order to learn video frame consistency for capturing temporal artifacts, we design a detection method based on the Consistency of video Frames (DeCoF), in which the core verification module consists of two transformer layers and an MLP head as shown in Figure 8. More specifically, given a video input $V$, the frozen parameter image encoder (ViT-L/14) extracts the features of the video frames and generates a feature sequence $S=\{M(F_{1}),M(F_{2}),\cdots,M(F_{L})\}$ with a shape of $L*D$, where $D$ represents the dimension of each feature. Next, the position encoding and a learnable class embedding representing the classification output are prepended to the sequence $S$ and then passed through a sequential module composed of two layers of transformers. Finally, a classification header is appended to the transformer layer, outputting a probability distribution over the target class. We train our model via a simple cross-entropy loss.
In addition to decoupling temporal and spatial artifacts, our approach significantly reduces computational complexity and memory requirements since we only need to learn anomalies between video frame features. As such, we achieve a simple yet effective approach.
6 Experiments
6.1 Experimental Setup
Data pre-processing. We only train on a certain sub-dataset of GVF and test on the remaining sub-datasets to evaluate the generalizability of the detector. We take first 8 frames of each video and center-crop each frame according to its shortest side, then resize it to 224 × 224. During the training process, some data augmentation, such as Jpeg compression, random horizontal flipping, and Gaussian blur are also used.
Evaluation metrics. Following previous methods [47, 48, 13], we mainly report accuracy (ACC) and average precision (AP) in our experiments to evaluate the detectors. The threshold for computing accuracy is set to 0.5 following [47, 48]. For image-based methods, we average frame-level predictions into corresponding video-level predictions.
Baseline. (i) CNNDet [47] detects images generated by GAN. Through a large amount of data enhancement, the model has generalizability on images generated by various GANs. (ii) DIF [39] is a few-shot learning method for image synthesis. (iii) DIRE [48] proposes a new image representation method called Diffusion Reconstruction Error (DIRE), which measures the error between the input image and its reconstruction via a pre-trained diffusion model to detect images generated by diffusion models. (iv) Lgard [43] uses a pre-trained CNN model as a transformation model to convert images into gradients and used these gradients to present generalized artifacts. (v) I3D [5] introduces a new Inflated 3D ConvNet that is based on 2D ConvNet inflation. (vi) Slow [12] proposes a dual-path model, but we only adopted the Slow Pathway, which is designed to capture the semantic information provided by a small number of sparse frames. (vii) X3D [11] designs a smaller and lighter structure and expanded it in different dimensions to achieve a trade-off between efficiency and accuracy. (viii) Mvit [10] combines the basic idea of multiscale feature hierarchies with transformer models for video and image recognition. (ix) UFDetect(UniversalFakeDetect) [29] detects generated images through pretrained CLIP:ViT visual encoders and nearest neighbors. (x) EVR [24] uses the CLIP backbone with frozen parameters and multi layer spatiotemporal Transformer decoder for video action recognition.
6.2 Evaluation
We conduct comprehensive evaluations to answer the following questions:
Table 2: Comprehensive comparison of our method with previous detectors and STCNN, including CNNDet, DIF, DIRE, Lgard, I3D, Slow, X3D and Mvit, were tested on our GVF dataset using the officially provided model. * Denotes we retrained it on a subdataset of GVF dataset with official code. We report ACC (%) and AP (%) (ACC and AP in the Table).
Method Training dataset Generation mode Detection Level Video generation models Total Avg.
Text2Video-zero Show1 ModelScope ZeroScope
ACC AP ACC AP ACC AP ACC AP ACC AP
CNNDet(0.5) - ProGAN Image 49.61 41.47 49.74 47.58 49.61 43.40 49.94 59.85 49.73 48.08
CNNDet(0.1) - ProGAN Image 49.93 46.38 50.06 46.72 50.19 42.97 50.06 55.27 50.06 47.84
DIF - Diffusion Image 35.10 47.96 45.90 53.42 31.40 46.91 27.50 46.59 34.98 48.72
DIRE - Diffusion Image 63.14 63.45 51.80 51.70 51.48 51.03 46.71 42.96 53.28 52.29
Lgard - ProGAN Image 53.09 64.78 59.70 68.42 40.21 38.87 47.94 51.80 50.24 55.97
UFDetect - ProGAN Image 50.13 63.99 53.09 74.42 52.13 75.55 58.89 84.27 53.56 74.56
CNNDet (0.5)* Text2Video-zero T2V Image 98.64 99.97 51.93 61.64 51.54 68.91 54.96 72.51 64.27 75.76
CNNDet (0.1)* Text2Video-zero T2V Image 96.84 99.99 53.93 60.99 55.34 68.87 60.69 73.80 66.70 75.91
DIRE* Text2Video-zero T2V Image 99.17 99.99 49.93 56.74 52.83 77.58 51.28 67.12 63.30 75.36
Lgard* Text2Video-zero T2V Image 98.45 100.00 56.19 78.91 51.03 66.30 54.64 75.93 65.08 80.29
I3D* Text2Video-zero T2V Video 98.00 99.87 49.75 70.41 49.25 71.43 50.50 51.85 61.88 73.39
Slow* Text2Video-zero T2V Video 93.30 99.48 50.48 52.09 50.59 49.43 49.41 37.29 60.95 59.57
X3D* Text2Video-zero T2V Video 94.84 98.61 49.41 33.52 48.35 43.07 49.85 30.89 60.61 51.52
Mvit* Text2Video-zero T2V Video 98.42 99.97 49.96 45.93 49.49 38.98 51.68 32.60 62.39 54.37
EVR* Text3Video-zero T2V Video 97.50 99.93 78.5 90.96 59.75 77.97 64.75 85.46 67.67 88.58
DeCoF Text2Video-zero T2V Video 99.50 100.00 71.50 95.83 78.25 96.13 76.50 96.13 81.44 97.02
DeCoF Show1 T2V Video 98.50 99.77 97.00 99.33 90.25 97.96 93.75 99.08 94.88 99.04
DeCoF ModelScope T2V Video 98.50 99.94 80.25 95.68 97.00 99.53 93.75 98.93 92.38 98.52
DeCoF ZeroScope T2V Video 97.00 99.81 75.50 96.52 75.75 96.42 98.50 99.91 86.69 98.17
Q1. How does our detector generalize to unseen T2V models compared to previous detectors?
Table 2 compares CNNDet [47], DIF [39], DIRE [48], Lgard [43], UFDetect [29], using the pre-trained weights provided by the official repository. Although these detectors have good performance in detecting generated images, their performance drops significantly on the task of detect generated videos, with both ACC and AP below 60%. As a result of retraining the above five detectors on the sub-dataset Text2Video-zero, their performance on seen video generation models greatly improved, but they still performed poorly for unseen generation models, with ACC and AP scores still lower than 60%, and the retrained STCNN presents similar results. On the contrary, our DeCoF has excellent generalizability for various unseen generation models. Compared with previous video classification methods based on pretrained CLIP [24], our method has still made significant progress. Specifically, on the task of generated video detection, DeCoF outperforms previous detectors by 13.77% $\sim$ 31.38% in ACC and 8.44% $\sim$ 49.18% in AP.
We also conducted cross-dataset evaluations. As shown in Table 2, the ACC of the model trained on the other three data sets is much higher than that of the model trained on Text2Video-zero by 5.25% $\sim$ 13.44%.
Figure 9: Robustness to unseen perturbations. The upper row represents the robustness of Gaussian blur, and the lower row represents the robustness of JPEG compression. We test the detector trained on different sub-datasets (represented by different colors) on different sub-datasets and use the AP (%) report for robustness comparison.
Q2. How does our detector maintain robustness against noise and disturbances?
In practical detection scenarios, the robustness of the detector to unseen perturbations is also crucial. Here, we mainly focus on the impact of two perturbations on the detector: Gaussian blur and JPEG compression. The perturbations are added under three levels for Gaussian blur ($\sigma$ = 1, 2, 3) and five levels for JPEG compression (quality = 90, 80, 70, 60, 50). Figure 9 shows the performance of our model trained on different sub-datasets in the face of the above two perturbations. Although experiencing some performance inconsistencies due to differences between training datasets, our method remains effective under all degrees of arbitrary perturbation, with the AP higher than 80%.
Q3. How does our detector perform on commercial generation models?
Table 3: Performance of DeCoF on commercial generation models.
Training dataset Video generation models
Gen-2 Pika Sora Veo
ACC AP ACC AP ACC AP ACC AP
Text2Video-zero 99.49 100.00 99.49 99.98 96.87 99.70 96.43 100.00
Show1 98.46 99.96 98.46 99.85 94.79 98.26 92.86 100.00
ModelScope 98.97 99.99 96.89 99.73 94.79 98.83 89.29 98.47
ZeroScope 98.97 99.99 94.84 99.52 89.58 97.92 67.86 94.90
To further evaluate the effectiveness of our detection method, we constructed four additional test datasets using well-known commercial proprietary T2V models, Pika [32], Gen-2 [15], Sora [41] and Veo [45]. As shown in Table 3, the detector achieved at least 94.84% ACC and 99.52% AP on Pika and at least 98.46% ACC and 99.96% AP on Gen-2. Due to regional restrictions, we did not qualify for testing Sora and Veo directly, but we constructed test sets for 48 public videos released by Sora and 14 public videos released by Veo, respectively. In this test, we achieved no less than 89.58% ACC and no less than 97.92% AP on Sora, no less than 67.86% ACC and no less than 94.90% AP on Veo, demonstrating the strong generalizability of our model DeCoF for unseen commercial T2V models.
7 Conclusions and Future Work
The paper presents the first investigation on generated video detection, with the goal of tackling the societal concern of validating the authenticity of video content. We construct the first public dataset for video detection, providing a solid benchmark for community research. At the same time, we propose a simple yet effective detection model, and prove its strong generalizability and robustness on multiple open-source and commercial closed-source models. We hope this study will inspire the creation and improvement of other detection technologies, providing new avenues for the development of authentic and reliable AIGC applications.
Appendix A Appendix
This supplementary material provides:
  • Sec. B: A detailed description of dataset construction, including dataset structure diagrams, as well as specific examples of prompts and videos.
  • Sec. C : Various parameter details for model training, hyperparameter settings, etc.
  • Sec. D : The impact of selecting different mapping functions.
  • Sec. E : The impact of video frame count.
  • Sec. F : The impact of video frame sampling interval.
  • Sec. G : More details on Sora.
  • Sec. H : Spatial artifacts vs Temporal artifacts.
  • Sec. I : Impact Statements.
  • Sec. J : Limitation.
  • Sec. K : More visual examples.
Appendix B $GeneratedVideoForensics$ Dataset
B.1 Dataset Composition
The following section provides a more detailed description of how the $GeneratedVideoForensics$ was constructed and evaluated. The entire GVF dataset is generated in two steps:
1) Selection of prompts and real videos: The prompts and real videos of the GVF dataset come from existing open-domain text-to-video datasets: MSVD  [6] and MSR-VTT  [50] datasets. When selecting samples, we followed the work of FETV  [25], Considering the collection of prompts from the above four aspects to simulate the sample distribution in real scenarios. More specifically, in the work of FETV, they classify each prompt from two aspects: the main content it describes and the attributes it controls during video generation. The spatial major content of the prompt can be divided into people, animals, plants, food, vehicles, buildings, artifacts, scenery, and illustrations. The corresponding spatial attribute controls are divided into quantity, color, and camera view. Likewise, the major contents and attribute controls of temporal are actions, kinetic motions, light change, and fluid motions, respectively, motion direction, speed, and event sequence.
2) Selection of generation models: Although there are many high-quality generation models on the market, we do not want to choose the best. On the contrary, we have selected four generation models with their advantages, aiming to simulate the crises we are encountering in real life: a large number of generation models appear, how do we deal with the problems caused by unseen generation models? In the end, the four most representative generation models were selected as our generation models: Text2Video-zero [22], ModelScopeT2V [46], ZeroScope[42], Show1 [51]. At the same time, we also selected the three most popular models as the secret test dataset: Pika [32], Gen-2 [15], Sora[41] and Veo [45]. For Sora and Veo, we provide detailed information in the section G
If divided by content, the $GenerateVideoForensics$ data set $D$ consists of 964 triples, which can be expressed as $D=\{(R,P,F)^{(i)}\}_{i=1}^{964}$, $R$ and $P$ represent real videos and corresponding text prompts, $F$ represents multiple fake videos generated by different generation models based on text prompts $P$.
If divided by the generation model, the GenerateVideoForensics data set can be divided into four sub-datasets: Text2Video-zero, ModelScopeT2V, ZeroScope, and Show-1. Each sub-dataset consists of 964 new triples, which can be expressed as $D=\{(R,P,f)^{(i)}\}_{i=1}^{964}$, $R$ and $P$ represent the real video and the corresponding text prompt, and $f$ represents the fake video generated by the specific generation model represented by the name of the current sub-dataset based on the text prompt $P$.
B.2 Dataset Evaluation
As part of this research, we present a comprehensive analysis of the GVF dataset, which includes an analysis of the attribute statistics of prompts, as well as an assessment of the different generation models in terms of the spatial content, spatial attributes, temporal content, and temporal attributes according to the FETV [25].
Table 4: Some examples of prompts. Each prompt is composed of a random combination of four attributes: main content of time, main attributes of time, main content of space, and main attributes of space. Here we list some examples of prompts, red for content and blue for attributes.
Main content Attribute
control
Example
Spatial People Quantity four friends are driving in the car
Animals Color a white bird chirps and plays on a handmade staircase
Plants Camera View a beautiful view of greenery and river from a top view
Food Color in cup full of chocolate then we add yellow cream to it so yummy
Temporal Actions Event Order a red-headed woman is shown walking and then sitting
Actions Speed a man fastly cross the road by run and get in to the car
Kinetic Motions Motion Direction the big black tractor is moving away from it s carrier
Fluid Motions Speed a bucket of slime is quickly poured on top of a woman’s head
Attribute statistics of prompts. From the four aspects mentioned before, the main content of space and time and the main attributes of space and time, we provide attribute statistics of prompts of the data set. As the figure 10 shows, the various attributes of prompts are different because one prompt may have multiple attributes at the same time. At the same time, different attributes have different difficulties in generating videos, and the harm they cause in a trust crisis is also different. Here, we also provide some specific examples of prompts, As shown in the table 4.
Figure 10: Data distribution over categories under the “major content” (upper) and “attribute control” (lower) aspects.
Evaluation of different generation models. Here, we present our assessment of four generation models according to the FETV standard [25]. We assess the generated videos based on four perspectives: (i) Static quality focuses on the visual quality of single video frames. (ii) Temporal quality focuses on the temporal coherence of video frames. (iii) Overall alignment measures the overall alignment between a video and the given text prompt. (iv) Fine-grained alignment measures the video-text alignment regarding specific attributes.
As shown in figure 11, Regarding spatial categories, the performance quality of these four data sets varies, and the performance of the same data set differs significantly between the static quality and temporal quality levels. The videos that include people, animals, and vehicles have poor spatial quality, but the videos that include plants, scenery, and natural items have greater spatial quality. The reason behind this is that the videos in the first three categories contain numerous high-frequency intricacies, such as face and finger, which are more difficult to comprehend compared to the low-frequency information found in the final two categories, such as water and clouds.
In terms of static quality, Text2Video-zero performs well but performs poorly in terms of temporal quality. Text2Video-zero inherited the ability to generate images from Stable Diffusion, but it lacks temporal coherence in its videos due to the lack of video training. Show-1 is second only to Text2Video-zero in terms of static quality, and its temporal quality is also superior. In comparison to ModelScopeT2V, ZeroScope is inferior to ModelScope in temporal quality but performs worse in static quality. The reason for this may be that ZeroScope has been primarily designed for videos with a 16:9 aspect ratio and without watermarks, rather than focusing on improving visual quality and aligning video text.
It is important to note that while the four models have better control over color and camera view, they do not perform as well when it comes to accurately controlling attributes such as quantity, motion direction, and order of the events. In terms of speed and motion direction attribute control, ModelScope and ZeroScope are particularly promising, whereas Text2Video-zero demonstrates poor performance.
On different attributes of the four models, fine alignment and global alignment perform similarly, although there are some exceptions. This is due to the fact that other properties may have an effect on the overall alignment. Therefore, it is more accurate to assess controllability using fine alignment when examining a specific attribute, whereas global alignment is a reasonable approximation.
Figure 11: The manual evaluation provided results across major static and temporal contents and attributes to control video-text alignment.
Appendix C Parameter Description
In this section, we provide a detailed description of various parameters of the experiment, including the hyperparameters of the model and the details of its training.
Training details. During the training process, we only train on one sub-dataset of the GVF dataset and test on the other sub-datasets. The numbers of the training set, validation set, and test set are 771, 96, and 97, respectively. We freeze the parameters of CLIP:ViT [8], making it a pure mapping function, and only train our model using the SGD optimizer with an initial learning rate of 0.001 and a momentum coefficient of 0.9. All experiments were completed on two RTX 3090 graphics cards.
Data pre-processing. All the following experiments are performed on our GVF dataset. In order to better detect the videos generated by the unseen generate model, we only train the model on the sub-dataset text2video-zero and test on the remaining three sub-datasets Show1 [51], ModelScope [46] and ZeroScope [42] Since the resolution and length of the videos generated by the four generation models are inconsistent, We take eight frames of each video for training and testing and center-crop each frame according to its shortest side, then resize it to 224×224. During the training process, we first resized the image to 256×256 according to bilinear interpolation and then randomly cropped with the size of 224 × 224. Some data augmentation, such as Jpeg compression, random horizontal flipping, and Gaussian blur, are also used.
Model hyperparameters. Our model consists of two layers of transformer and an MLP head is an adaptation of the sequence to sequence trans- former, where the dimension and feature dimension of MLP is 768, path_size is 32, the dropout parameter is 0.1, and the number of attention heads is 4.
Appendix D Different Mapping Functions
In this section, we examine the impact of different mapping functions.
The mapping function should have two properties. One is that the distance between features in the feature space is inversely proportional to the similarity between images, and the other is that it is insensitive to spatial artifacts. The second requirement is easy to implement as long as the mapping function is not used to train the fake content detection task. For the first requirement, we explored several different mapping functions, as shown in the table 5.
First, we discussed the two different tasks of ImageNet and CLIP. Compared to ImageNet’s image classification task, CLIP’s image encoder trained on text image task alignment is a better mapping function. We considered two reasons: one is the size of the training data, and the other is due to the difference in the task itself; We do not want the model to classify images into specific items or biological categories but rather cluster similar images. At the same time, we compared the impact of different model frameworks: Due to issues such as the number of model parameters, Resnet trained on the CLIP task is not suitable as a mapping function.
Overall, the mapping function largely determines the effectiveness of our approach.
Table 5: A comparison of the impact of different mapping functions.
Mapping Function Pretrain Tasks Video generation models Total
Avg.
Text2Video-zero Show1 ModelScope ZeroScope
ACC AP ACC AP ACC AP ACC AP ACC AP
ViT-L/14 CLIP 99.50 100.00 71.50 95.83 78.25 96.13 76.50 96.13 81.44 97.02
Resnet50 CLIP 96.92 99.82 52.60 80.66 57.22 80.19 50.00 67.65 64.19 82.08
ViT-B/16 ImageNet 98.00 99.79 65.39 87.65 54.70 76.39 55.57 76.1 68.41 84.98
Resnet50 ImageNet 99.5 99.96 54.2 81.9 51.43 67.97 49.93 58.65 63.76 77.12
Appendix E The Impact of Video Frame Number
Here we discuss the impact of the number of video frames used for detection.
Table 6: A comparison of the impact of different frame numbers.
Training dataset Frame Numbers Video generation models Total
Avg.
Text2Video-zero Show1 ModelScope ZeroScope
ACC AP ACC AP ACC AP ACC AP ACC AP
Text2Video-zero 8 99.50 100.00 71.50 95.83 78.25 96.13 76.50 96.13 81.44 97.02
Show1 8 98.50 99.77 97.00 99.33 90.25 97.96 93.75 99.08 94.88 99.04
ModelScope 8 98.50 99.94 80.25 95.68 97.00 99.53 93.75 98.93 92.38 98.52
ZeroScope 8 97.00 99.81 75.50 96.52 75.75 96.42 98.50 99.91 86.69 98.17
Text2Video-zero 16 99.50 99.99 60.70 95.13 65.55 93.55 63.27 95.38 72.25 96.01
Show1 16 97.86 99.69 95.93 99.04 93.73 98.13 94.30 98.88 95.45 98.93
ModelScope 16 100.00 100.00 81.25 95.41 98.36 99.55 95.36 99.19 93.74 98.53
ZeroScope 16 97.43 99.83 80.75 96.10 82.18 96.26 96.93 99.72 89.32 97.97
Text2Video-zero 24 99.50 100.00 57.57 95.14 55.7 94.10 60.18 96.32 68.23 96.39
Show1 24 97.80 99.98 94.36 99.10 89.66 97.49 93.73 98.93 93.88 98.87
ModelScope 24 100.00 100.00 74.68 95.57 97.00 99.32 90.73 99.30 90.60 98.54
ZeroScope 24 98.43 99.96 83.32 97.00 85.02 96.56 97.50 99.78 91.06 98.32
In our work, we selected 8 frames for generated video detection for two reasons: on the one hand, the current generation model still has this deficiency in generating long videos. Even the videos currently displayed by Sora can only generate about 300-600 frames, so we believe that 8 frames is an acceptable choice. On the other hand, in order to highlight the effectiveness of our method, we chose a challenging number of 8 frames, and a large number of experiments have also proven that our method indeed performs very well.
We conducted experiments on the impact of frames 8, 16, and 24 on detection, as shown in the table 6. We found that the performance of 16 frames is generally better than that of 8 frames, except for the models trained on the Text2Video zero subdataset, which is consistent with our previous evaluation results of several generation models. Text2Video zero has the worst time continuity. In addition, the effect will become worse at 24 frames. One possible reason is that as the sequence increases, the temporal artifacts of the generated video become more unique, which weakens the model’s generalizability.
Appendix F The Impact of Video Frame Sampling Interval.
In this section, we discussed the impact of sampling interval on detection results. In previous experiments, we usually took the first 8 frames of the video for detection.
Unlike previous experiments, this experiment was conducted on the latest generation model Sora, as Sora has made significant progress in generating frames compared to other generation models, making it easier for us to comprehensively discuss the impact of sampling strategies. As shown in the table 7, facing three different sampling strategies: 1) Take the first 8 frames; 2) Take 8 frames evenly from all frames; 3) Take 8 frames evenly from the first 300 frames; Our method did not show a significant decrease, which we attribute to two reasons: 1) The prompts we collected were comprehensive enough to cover the speed of video changes well. 2) The duration of the test video is still around 10 seconds, so the impact of sampling interval is not very significant. For longer generated videos, the performance degradation caused by changes in sampling intervals will be more severe.
Table 7: The impact of different sampling strategies
Training dataset Sampling strategy Sora
ACC AP
Text2Video-zero First 8 frames 96.87 99.70
Show1 First 8 frames 94.79 98.26
ModelScope First 8 frames 94.79 98.83
ZeroScope First 8 frames 89.58 97.92
Text2Video-zero Take 8 frames evenly from all frames 96.87 99.48
Show1 Take 8 frames evenly from all frames 94.79 98.61
ModelScope Take 8 frames evenly from all frames 94.79 98.96
ZeroScope Take 8 frames evenly from all frames 88.54 98.31
Text2Video-zero Take 8 frames evenly from first 300 frames 97.91 99.65
Show1 Take 8 frames evenly from first 300 frames 94.79 98.48
ModelScope Take 8 frames evenly from first 300 frames 93.75 98.74
ZeroScope Take 8 frames evenly from first 300 frames 90.62 98.22
Appendix G More Details on Sora.
G.1 Test data on Sora
Due to regional restrictions, we were not eligible to test Sora. But we tested all of the 48 videos officially displayed by Sora. The test dataset consists of 48 official videos released by Sora and 48 real videos from the original test set, the test on Veo are similar.
G.2 The detection performance of human testers on Sora
We invited three testers to evaluate the authenticity of 96 videos (48 real and 48 Sora generated) from the previously mentioned Sora test set. The experimental results are shown in the table 8.
Table 8: Compared to the test results of human testers
Training dataset Sora
TP+TN ACC
Tester 1 57(48+9) 59.37
Tester 2 66(47+19) 68.75
Tester 3 60(48+12) 62.5
Text2Video-zero 93 96.87
Show1 91 94.79
ModelScope 91 94.79
ZeroScope 86 89.58
The results of DeCoF in testing far exceed the level of human judgment. At the same time, we also need to point out that interviews with testers after testing revealed that their judgment on the videos generated by Sora was artificially high, as there were 4-5 flashy videos in the officially released videos that did not match the results that could appear in the real world, such as pirate ships in coffee cups. Meanwhile, according to the testers, their judgment is based on whether there is irregular movement (such as objects appearing out of thin air) or object threading (such as overlapping basketball and basket) and abnormal distortion (such as human fingers or limbs of moving animals) in the generated video.
Appendix H Spatial artifacts vs Temporal artifacts.
H.1 Detect generated videos from temporal artifacts
Although our method attempts to detect generated videos from temporal artifacts, there is a question: is it possible that CLIP: ViT has differences in the features between the generated video and the real video frames when extracting image features, resulting in better detection performance? In other words, is our method really detect generated videos from temporal artifacts?
To prove this, we repeated the previous probe experiment. As shown in the table 9, our model has also performed well in detecting test data without spatial artifacts. On the other hand, for the second set of test data in the previous probe experiment, our model has also successfully detected it. We believe that repeated stillness in multiple frames is also a type of temporal artifact to some extent.
Table 9: The proof of our method detect generated videos from temporal artifact
Method Video generation models
Temporal Spatial
ACC AP ACC AP
I3D 51.17 51.61 86.67 98.61
Slow 50.04 50.37 93.81 99.51
X3D 50.04 54.31 78.39 92.84
Mvit 50.48 55.29 72.16 92.59
Ours 98.46 99.80 98.46 99.66
H.2 Spatial artifacts vs Temporal artifacts.
In this paper, we attempt to detect generated video from the perspective of temporal artifacts, but this does not mean that it is impossible to detect generated video from spatial artifacts. In fact, the difficulty of detecting spatial artifacts in generated videos lies in how to extract more general artifacts, as different generation models generate videos with different resolutions, color brightness, contrast, and so on.
Appendix I Impact Statements
Our research focuses on detecting generated videos. We developed the DeCoF model and shared our dataset to support detecting generated videos. These works are crucial to protect digital content and prevent the spread of disinformation. Nevertheless, these tools have the potential to be misused, resulting in a rivalry between video generation and detection technology. We want to advocate for the ethical use of technology and foster research in the creation of tools for verifying media authenticity. We are confident that this will result in protecting the public from the harms of disinformation, improving the clarity and authenticity of information distribution, and ensuring the protection of personal privacy for individuals.
Appendix J Limitation
In this paper, we tackle an overlooked challenge, i.e., generated video detection, by proposing a universal detector to deal with the trust crisis caused by various emerging text-to-video (T2V) models, especially unseen models. However, there are various methods of video forgery, such as Deepfake and malicious editing, in which our methods may experience significant performance degradation or inapplicability. Taking Deepfake detection as an example, our method does not add additional detection modules to the face. In addition, Deepfake videos are often locally tampered with, and the generated videos are generated as a whole. Our work focuses on detecting generated videos, which has a more diverse range of forgery scenarios and more flexible forgery techniques.
Appendix K More Visual Examples
Here, we provide more examples of generated videos. Figures 12 to Figures 15 are examples of real videos and generated videos corresponding to prompts.
Figure 12: Real videos and videos generated by four generation models according to prompt: A person mixing ingredients in a mixing bowl.
Figure 13: Real videos and videos generated by four generation models according to prompt: A kitten moves about.
Figure 14: Real videos and videos generated by four generation models according to prompt: A young man is seated and playing a guitar.
Figure 15: Real videos and videos generated by four generation models according to prompt: A boy rides a bike on a fence.